Halliford Corporation expects to have earnings this coming year of $3.174 per share. Halliford plans to retain all of its earplease help!
Answer #1

Halliford’s dividend forecast (g = retention rate × return on new investment)

Year

0

1

2

3

4

5

6

Earnings

1).EPS Growth Rate (vs. prior yr)

23.80%

23.80%

11.186%

11.186%

4.284%

2). EPS

$3.174

$3.9294

$4.8646

$5.4088

$6.0138

$6.2714

Dividends

3). Retention Ratio

100%

100%

47%

47%

18%

18%

4). Dividend Payout Ratio

0%

0%

53%

53%

82%

82%

Dividends(2 x 4) $0 $0    $2.5782 $2.8666 $4.9313 $5.1426

From year 5 on, dividends grow at constant rate of 4.284%. Therefore,

P(5) = D6 / (r - gC) = $5.1426 / (10.5% - 4.284%) = $5.1426 / 0.06216 = $82.73

P(0) = [D3 / (1 + r)3] + [D4 / (1 + r)4] + [(D5 + P5) / (1 + r)5]

= [$2.5782 / 1.1053] + [$2.8666 / 1.1054] + [($4.9313 + $82.73) / 1.1055]

= $1.91 + $1.92 + $53.21 = $57.04